OpenStudy (anonymous):

give the exact value of f(pi/3) if f(X)=6cosx-10sin2x

OpenStudy (anonymous):

why 7x+4?

OpenStudy (aravindg):

substitute \(x\) with \(\dfrac{\pi}{3}\)

OpenStudy (anonymous):

ok and then from here I am not clear what do

OpenStudy (anonymous):

Do you know what is the cosine and sine of pi/3?

OpenStudy (anonymous):

cosine pi/3=.9998329795 and sine pi/3=.0182760276

OpenStudy (aravindg):

:) Then whats the problem ?

OpenStudy (anonymous):

Do you know what the unit circle is?

OpenStudy (aravindg):

just put those values in the function get the answer be happy :)

OpenStudy (anonymous):

The values are not correct. cosine of pi/3 is .5

OpenStudy (anonymous):

and sine of pi/3= sqrt(3)/2

OpenStudy (anonymous):

ok. I have the answer my teacher put up on the board which is 3-5\[\sqrt3\]

OpenStudy (anonymous):

Remember that you are finding the values in radians, not degrees.

OpenStudy (anonymous):

but I cannot seem to get this value

OpenStudy (anonymous):

oh! that helps thanks :)

OpenStudy (anonymous):



OpenStudy (anonymous):

Just continue from there.

OpenStudy (anonymous):



OpenStudy (anonymous):

oh ok I got it! thank you so much!

OpenStudy (anonymous):

Sure.